If two coins, A and B are tossed simultaneously, find the probability of one-coin landing heads up and the other landing tails up in the following order.

a) Probability of coin A landing heads up

b) Probability of coin B landing tails up

c) Probability of coin A landing heads up and coin B tails up

d) Probability of coin A landing tails up

e) Probability of coin A landing tails up and coin B landing heads up


1
Expert's answer
2021-11-09T16:52:32-0500

P(coin landing on head) = P(coin landing on tail) = 1/2 = 0.5

a) P(coin A landing heads up) = 0.5

b) P(coin B landing tails up) = 0.5

c) P(coin A landing heads up and coin B landing tails up) = P(coin A landing heads up) * P(coin B landing tails up) = 0.5 * 0.5 = 0.25

d) P(coin A landing tails up) = 0.5

e) P(coin A landing tails up and coin B landing heads up) = P(coin A landing tails up) * P(coin B landing heads up) = 0.5 * 0.5 = 0.25